# Multispecies inhomogeneous $t$-PushTASEP from antisymmetric fusion

### v2

@@ Appendix A $S^{k,1}(z)$ from antisymmetric fusion
Appendix B $H_{\text{PushTASEP}}(x_{1},\ldots,x_{L})$ from transfer matrices for symmetric fusion
@@
For comparison, we briefly sketch an alternative description of $\mathcal{H}$ (51) in terms of transfer matrices corresponding to symmetric fusion. We introduce the symmetric tensor counterparts of (13) and (14). For $k\in{\mathbb{Z}}_{\geq 0}$, define
$\mathscr{B}_{k}$ =$=\{{\bf i}=(i_{0},\ldots,i_{n})\in({\mathbb{Z}}_{\geq 0})^{n+1}\mid|{\bf i}|=k\},\quad(|{\bf i}|=i_{0}+\cdots+i_{n}),$ (108)
$V_{k}$ $=\bigoplus_{{\bf i}\in\mathscr{B}_{k}}\mathbb{C}v^{\prime}_{\bf i}.$ (109)
@@
In the special case $k=1$, we identify $V_{1}$ with ${\mathsf{V}}$, the space of local states of the $t$-PushTASEP defined in Section 2, following the same rule as (15), with $v^{\prime}_{{\bf e}_{i}}\in V_{1}$ replacing $v^{\prime}_{{\bf e}_{i}}\in V^{1}$. That is,
@@
${\mathsf{V}}\ni{\mathsf{v}}_{i}=v^{\prime}_{{\bf e}_{i}}\in V_{1}\;\;\text{where}\;\;{\bf e}_{i}=(\delta_{0,i},\ldots,\delta_{n,i})\in\mathscr{B}_{1}\quad(0\leq i\leq n).$ (110)
@@
A quantum $R$-matrix acts on $V_{k}\otimes V_{1}$ and satisfies the Yang-Baxter equation. To distinguish it from $S(z)=S^{k,1}(z)$ in (29)–(33), we use a a different notation: $\mathscr{S}(z)=\mathscr{S}_{k,1}(z)$. Employing a suitable gauge (cf. [KOS24, eq. (86)]), we have
$\mathscr{S}(z)(v^{\prime}_{\bf i}\otimes v^{\prime}_{{\bf e}_{j}})$ $=\sum_{{\bf a}\in\mathscr{B}_{k},{\bf e}_{b}\in\mathscr{B}_{1}}\mathscr{S}(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}\,v^{\prime}_{\bf a}\otimes v^{\prime}_{{\bf e}_{b}}\qquad({\bf i}\in\mathscr{B}_{k},{\bf e}_{j}\in\mathscr{B}_{1}),$ (111)
$\mathscr{S}(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}$ $=\delta^{{\bf a}+{\bf e}_{b}}_{{\bf i}+{\bf e}_{j}}t^{i_{b+1}+\cdots+i_{n}}(1-t^{i_{b}}z^{\delta_{b,j}})z^{[j>b]}.$ (112)
@@
The most significant difference from (33) is that (112) is defined for ${\bf a},{\bf i}\in\mathscr{B}_{k}$ (108) rather than $\mathscr{B}^{k}$ (13).
@@
For $k\in{\mathbb{Z}}_{\geq 0}$, let $T_{k}(z)=T_{k}(z|x_{1},\ldots,x_{L})$ be the transfer matrix whose auxiliary space is $V_{k}$. Following the construction in (36a)– (36b), it is given by
$T_{k}(z)|\sigma_{1},\ldots,\sigma_{L}\rangle$ $=\sum_{\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}\in\{0,\ldots,n\}}T_{k}(z)^{\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}}_{\sigma_{1},\ldots,\sigma_{L}}|\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}\rangle,$ (113a)
@@
$T_{k}(z)^{\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}}_{\sigma_{1},\ldots,\sigma_{L}}$ $=\sum_{{\bf a}_{1},\ldots,{\bf a}_{L}\in\mathscr{B}_{k}}\mathscr{S}\Bigl{(}\frac{z}{x_{1}}\Bigr{)}^{{\bf a}_{2},{\bf e}_{\sigma^{\prime}_{1}}}_{{\bf a}_{1},{\bf e}_{\sigma_{1}}}\mathscr{S}\Bigl{(}\frac{z}{x_{2}}\Bigr{)}^{{\bf a}_{3},{\bf e}_{\sigma^{\prime}_{2}}}_{{\bf a}_{2},{\bf e}_{\sigma_{2}}}\cdots\mathscr{S}\Bigl{(}\frac{z}{x_{L}}\Bigr{)}^{{\bf a}_{1},{\bf e}_{\sigma^{\prime}_{L}}}_{{\bf a}_{L},{\bf e}_{\sigma_{L}}}.$ (113b)
@@
$[T_{k}(z|x_{1},\ldots,x_{L}),T_{k^{\prime}}(z^{\prime}|x_{1},\ldots,x_{L})]$ $=0\qquad(k,k^{\prime}\in{\mathbb{Z}}_{\geq 0}),$ (114a)
$[T_{k}(z|x_{1},\ldots,x_{L}),T^{l}(z^{\prime}|x_{1},\ldots,x_{L})]$ $=0\qquad(k\in{\mathbb{Z}}_{\geq 0},l\in\{0,\ldots,n+1\})$ (114b)
in addition to (37).
@@
$T_{l}(z)$ $=\left(T^{0}(tz)T^{0}(t^{2}z)\cdots T^{0}(t^{l-1}z)\right)^{-1}\mathrm{det}\left(T^{1-i+j}(t^{j-1}z)\right)_{1\leq i,j\leq l},$ (115a)
$T^{k}(z)$ $=\left(T^{0}(t^{-1}z)T^{0}(t^{-2}z)\cdots T^{0}(t^{-k+1}z)\right)^{-1}\mathrm{det}\left(T_{1-i+j}(t^{1-j}z)\right)_{1\leq i,j\leq k},$ (115b)
where $T^{0}(z)$ is a scalar matrix (39), and we set $T^{k}(z)=0$ for $k<0$ or $k>n+1$. Substituting (115b) into $\mathcal{H}$ in (51) provides an alternative expression for $H_{\text{PushTASEP}}$ in terms of differential coefficients of $T_{0}(z),T_{1}(z),\ldots,T_{n}(z)$. However the resulting formula is not particularly illuminating. For instance for $n=2$, we obtain
$\mathcal{H}$ $=\frac{\dot{T}_{2}(0)-(1+t^{m_{0}}+t^{1+m_{0}}+t^{m_{0}+m_{1}}+t^{1+m_{0}+m_{1}})\dot{T}_{1}(0)+tC\sum_{j=1}^{L}\frac{1}{x_{j}}}{(1-t)t(1-t^{m_{0}})(1-t^{m_{0}+m_{1}})},$ (116a)
$C$ $=-1+t-t^{-1+m_{0}}-t^{2m_{0}}-t^{1+m_{0}}-t^{-1+m_{0}+m_{1}}-t^{2(m_{0}+m_{1})}-t^{1+m_{0}+m_{1}}-t^{-1+2m_{0}+m_{1}}-2t^{2m_{0}+m_{1}},$ (116b)

### v3

@@ B.1. Modifying $R(z)$ into $S(z)$
Appendix C $H_{\text{PushTASEP}}(x_{1},\ldots,x_{L})$ from transfer matrices for symmetric fusion
@@
For comparison, we briefly sketch an alternative description of $\mathcal{H}$ (5.1) in terms of transfer matrices corresponding to symmetric fusion. We introduce the symmetric tensor counterparts of (3.1) and (3.2). For $k\in{\mathbb{Z}}_{\geq 0}$, define
$\mathscr{B}_{k}$ $=\{{\bf i}=(i_{0},\ldots,i_{n})\in({\mathbb{Z}}_{\geq 0})^{n+1}\mid|{\bf i}|=k\},\quad(|{\bf i}|=i_{0}+\cdots+i_{n}),$ (C.1)
$V_{k}$ $=\bigoplus_{{\bf i}\in\mathscr{B}_{k}}\mathbb{C}v^{\prime}_{\bf i}.$ (C.2)
@@
In the special case $k=1$, we identify $V_{1}$ with $\mathsf{V}$}, the space of local states of the $t$-PushTASEP defined in Section2, following the same rule as (3.3), with $v^{\prime}_{{\bf e}_{i}}\in V_{1}$ replacing $v^{\prime}_{{\bf e}_{i}}\in V^{1}$. That is,
@@
${\mathsf{V}}\ni{\mathsf{v}}_{i}=v^{\prime}_{{\bf e}_{i}}\in V_{1}\;\;\text{where}\;\;{\bf e}_{i}=(\delta_{0,i},\ldots,\delta_{n,i})\in\mathscr{B}_{1}\quad(0\leq i\leq n).$ (C.3)
@@
A quantum $R$-matrix acts on $V_{k}\otimes V_{1}$ and satisfies the Yang-Baxter equation. To distinguish it from $S(z)=S^{k,1}(z)$ in (B.13)–(B.17), we use a a different notation: $\mathscr{S}(z)=\mathscr{S}_{k,1}(z)$. Employing a suitable gauge (cf. [KOS24, eq. (86)]), we have
$\mathscr{S}(z)(v^{\prime}_{\bf i}\otimes v^{\prime}_{{\bf e}_{j}})$ $=\sum_{{\bf a}\in\mathscr{B}_{k},{\bf e}_{b}\in\mathscr{B}_{1}}\mathscr{S}(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}\,v^{\prime}_{\bf a}\otimes v^{\prime}_{{\bf e}_{b}}\qquad({\bf i}\in\mathscr{B}_{k},{\bf e}_{j}\in\mathscr{B}_{1}),$ (C.4)
$\mathscr{S}(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}$ $=\delta^{{\bf a}+{\bf e}_{b}}_{{\bf i}+{\bf e}_{j}}t^{i_{b+1}+\cdots+i_{n}}(1-t^{i_{b}}z^{\delta_{b,j}})z^{[j>b]}.$ (C.5)
@@
The most significant difference from (B.17) is that (C.5) is defined for ${\bf a},{\bf i}\in\mathscr{B}_{k}$ (C.1) rather than $\mathscr{B}^{k}$ (3.1).
@@
For $k\in{\mathbb{Z}}_{\geq 0}$, let $T_{k}(z)=T_{k}(z|x_{1},\ldots,x_{L})$ be the transfer matrix whose auxiliary space is $V_{k}$. Following the construction in (4.2a)– (4.2b), it is given by
$T_{k}(z)|\sigma_{1},\ldots,\sigma_{L}\rangle$ $=\sum_{\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}\in\{0,\ldots,n\}}T_{k}(z)^{\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}}_{\sigma_{1},\ldots,\sigma_{L}}|\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}\rangle,$ (C.6a)
@@
$T_{k}(z)^{\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}}_{\sigma_{1},\ldots,\sigma_{L}}$ $=\sum_{{\bf a}_{1},\ldots,{\bf a}_{L}\in\mathscr{B}_{k}}\mathscr{S}\Bigl(\frac{z}{x_{1}}\Bigr)^{{\bf a}_{2},{\bf e}_{\sigma^{\prime}_{1}}}_{{\bf a}_{1},{\bf e}_{\sigma_{1}}}\mathscr{S}\Bigl(\frac{z}{x_{2}}\Bigr)^{{\bf a}_{3},{\bf e}_{\sigma^{\prime}_{2}}}_{{\bf a}_{2},{\bf e}_{\sigma_{2}}}\cdots\mathscr{S}\Bigl(\frac{z}{x_{L}}\Bigr)^{{\bf a}_{1},{\bf e}_{\sigma^{\prime}_{L}}}_{{\bf a}_{L},{\bf e}_{\sigma_{L}}}.$ (C.6b)
@@
$[T_{k}(z|x_{1},\ldots,x_{L}),T_{k^{\prime}}(z^{\prime}|x_{1},\ldots,x_{L})]$ $=0\qquad(k,k^{\prime}\in{\mathbb{Z}}_{\geq 0}),$ (C.7a)
$[T_{k}(z|x_{1},\ldots,x_{L}),T^{l}(z^{\prime}|x_{1},\ldots,x_{L})]$ $=0\qquad(k\in{\mathbb{Z}}_{\geq 0},l\in\{0,\ldots,n+1\})$ (C.7b)
in addition to (4.3).
@@
$T_{l}(z)$ $=\left(T^{0}(tz)T^{0}(t^{2}z)\cdots T^{0}(t^{l-1}z)\right)^{-1}\mathrm{det}\left(T^{1-i+j}(t^{j-1}z)\right)_{1\leq i,j\leq l},$ (C.8a)
$T^{k}(z)$ $=\left(T^{0}(t^{-1}z)T^{0}(t^{-2}z)\cdots T^{0}(t^{-k+1}z)\right)^{-1}\mathrm{det}\left(T_{1-i+j}(t^{1-j}z)\right)_{1\leq i,j\leq k},$ (C.8b)
where $T^{0}(z)$ is a scalar matrix (4.5), and we set $T^{k}(z)=0$ for $k<0$ or $k>n+1$. Substituting (C.8b) into $\mathcal{H}$ in (5.1) provides an alternative expression for $H_{\text{PushTASEP}}$ in terms of differential coefficients of $T_{0}(z),T_{1}(z),\ldots,T_{n}(z)$. However the resulting formula is not particularly illuminating. For instance for $n=2$, we obtain
$\mathcal{H}=$ $\frac{\dot{T}_{2}(0)-(1+t^{m_{0}}+t^{1+m_{0}}+t^{m_{0}+m_{1}}+t^{1+m_{0}+m_{1}})\dot{T}_{1}(0)+tC\sum_{j=1}^{L}\frac{1}{x_{j}}}{(1-t)t(1-t^{m_{0}})(1-t^{m_{0}+m_{1}})},$ (C.9a)
$C=$ $-1+t-t^{-1+m_{0}}-t^{2m_{0}}-t^{1+m_{0}}-t^{-1+m_{0}+m_{1}}-t^{2(m_{0}+m_{1})}$ (C.9b)
$-t^{1+m_{0}+m_{1}}-t^{-1+2m_{0}+m_{1}}-2t^{2m_{0}+m_{1}},$ (C.9c)